# Variational Transformer Ansatz for the Density Operator of Steady States in Dissipative Quantum Many-Body Systems

### v1

@@ III.2 Mixed-State Observables
$\langle\hat{A}\rangle=\frac{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\,\hat{A}\bigr{)}}{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\bigr{)}}=\sum_{\bm{\alpha}\in\mathbf{M}}\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\bigr{)}}\,\tilde{A}_{\rho_{\bm{\theta}}}(\bm{\alpha}),$ (S26)
@@
$\tilde{A}_{\rho_{\bm{\theta}}}(\bm{\alpha})=\sum_{\bm{\beta}}\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}\langle\bm{\beta}|\hat{A}|\bm{\alpha}\rangle.$ (S27)
@@
plays the role of a classical distribution over the diagonal elements of $\hat{\rho}_{\bm{\theta}}$, and the local estimator $\tilde{A}_{\rho_{\bm{\theta}}}(\bm{\alpha})$ here involves an inner sum over all basis states (or a suitably chosen subset) to capture the off-diagonal contributions $\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})$ to the observable $\hat{A}$. In this manner, the quantum expectation value reduces to a standard classical average over a series of sampled configurations $\mathbf{M}$, allowing one to use the same Metropolis-Hastings sampling scheme described in Sec. III.1 to estimate both observables and their gradients for the mixed-state variational ansatz.
III.3 Metric Tensor in Stochastic Reconfiguration
@@
$S_{ij}=\langle\Delta O_{i}^{*}\Delta O_{j}\rangle-\langle\Delta O_{i}^{*}\rangle\langle\Delta O_{j}\rangle,$ (S28)
@@
$\Delta O_{i}=\frac{\partial\log\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}{\partial\theta_{i}}$ (S29)
@@
In the context of mixed-state variational ansatz, this metric tensor is computed using Monte Carlo sampling as described in Sec. III.2. Once $S$ is constructed, it is used in the natural gradient update step to precondition the gradient of the energy functional.
III.4 Regularization and Stabilization
@@
$S^{\prime}=S+\lambda\mathds{I},$ (S30)
@@
This is implemented as the stochastic reconfiguration method as a gradient preconditioner in NetKet, where $S$ is constructed from Monte Carlo estimates as described in Section III.3. This approach stabilizes the training dynamics of the transformer density operator ansatz and improves convergence in the steady-state optimization.
III.5 Optimization Procedure
@@
^$\hat{\mathcal{L}}|\rho_{\bm{\theta}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\hat{\mathcal{L}}\bm{\rho}_{\bm{\theta}}=0,$ (S31)
@@
$\operatorname{Loss}(\bm{\theta})=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho_{\bm{\theta}}|\mathfrak{L}|\rho_{\bm{\theta}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (S32)
@@
$\mathfrak{L}=\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}.$ (S33)
@@
$\operatorname{Cost}(\bm{\theta})=\frac{\|\hat{\mathcal{L}}\bm{\rho}_{\bm{\theta}}\|_{2}^{2}}{\|\bm{\rho}_{\bm{\theta}}\|_{2}^{2}}=\frac{\operatorname{Tr}\Bigl{(}\bm{\rho}_{\bm{\theta}}^{\dagger}\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}\bm{\rho}_{\bm{\theta}}\Bigr{)}}{\operatorname{Tr}\Bigl{(}\bm{\rho}_{\bm{\theta}}^{\dagger}{\bm{\rho}_{\bm{\theta}}}\Bigr{)}},$ (S34)
@@
In our implementation, we use a hybrid optimization approach that combines standard first-order gradient updates with second-order corrections via stochastic reconfiguration introduced in III.4. The Stochastic Gradient Descent or Adam optimizer performs standard gradient updates, while the stochastic reconfiguration accounts for the curvature of the variational manifold by introducing a metric tensor $S$, effectively implementing a natural gradient descent strategy. Their methods are provided in NetKet via a dedicated variational driver nk.SteadyState and Eq. (S34) is the cost function that NetKet uses in its steady-state driver.
III.5.1 Hybrid Optimization Approach
@@
${\bm{\theta}}^{(k+1)}={\bm{\theta}}^{(k)}-\eta\nabla_{\bm{\theta}}\operatorname{Cost}(\bm{\theta}),$ (S35)
@@
${\bm{\theta}}^{(k+1)}={\bm{\theta}}^{(k)}-\eta S^{-1}\nabla_{\bm{\theta}}\operatorname{Cost}(\bm{\theta}),$ (S36)
@@
The stochastic gradient $\nabla_{\bm{\theta}}\operatorname{Cost}(\bm{\theta})$ is estimated over the probability distribution defined by the entries of the vectorized density matrix $P_{\bm{\theta}}(\bm{\alpha},\bm{\beta})\propto|\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})|^{2}$ as in Sec.III.2. The gradient of the cost function with respect to the complex conjugate of the $i$th parameter can be expressed as
$\frac{\partial}{\partial\theta_{i}^{*}}\operatorname{Cost}(\bm{\theta})=\langle\tilde{\mathcal{L}}_{i}\nabla_{i}^{*}\tilde{\mathcal{L}}_{i}\rangle-\langle O_{i}^{*}\,\tilde{\mathcal{L}}^{2}\rangle,$ (S37)
@@
~$\tilde{\mathcal{L}}(\bm{\alpha},\bm{\beta})=\frac{\sum_{\bm{\alpha}^{\prime},\bm{\beta}^{\prime}}\hat{\mathcal{L}}(\bm{\alpha},\bm{\beta};\bm{\alpha}^{\prime},\bm{\beta}^{\prime})\,\rho_{\bm{\theta}}(\bm{\alpha}^{\prime},\bm{\beta}^{\prime})}{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}.$ (S38)
III.5.2 Optimization Workflow
@@
    Sample a batch of configurations from the current transformer density operator ansatz using a Markov chain Monte Carlo sampler introduced in Sec. III.1.
@@
    Compute the metric tensor $S$ for stochastic reconfiguration introduced in Sec. III.3.
@@
    Regularize the metric tensor with a small constant $\lambda$, i.e., $S^{\prime}=S+\lambda\,\mathds{I}$, to ensure numerical stability as introduced in Sec. III.4.
@@
III.5.3 Learning Rate Scheduling
@@
$\mathrm{lr}(i_{\mathrm{step}})=\begin{cases}\eta_{0},&i_{\mathrm{step}}<i_{\mathrm{switch}}\\ \eta_{0}\cdot\frac{1+\cos(\pi(i_{\mathrm{step}}-i_{\mathrm{switch}})/i_{\mathrm{decay}})}{2}+\alpha\eta_{0},&i_{\mathrm{step}}\geq i_{\mathrm{switch}}\end{cases}$ (S39)
@@
$\lambda_{\mathrm{SR}}(i_{\mathrm{step}})=\begin{cases}\lambda_{0},&i_{\mathrm{step}}<i_{\mathrm{switch,SR}}\\ \lambda_{0}\cdot\frac{1+\cos(\pi(i_{\mathrm{step}}-i_{\mathrm{switch,SR}})/i_{\mathrm{decay,SR}})}{2}+\alpha_{\mathrm{SR}}\lambda_{0},&i_{\mathrm{step}}\geq i_{\mathrm{switch,SR}}\end{cases}$ (S40)
@@
III.6 Benchmark Methods for Steady-State Computation
@@ III.6.1 Exact Diagonalization
$\mathfrak{L}=\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}.$ (S42)
@@
$\mathfrak{L}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\lambda|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (S43)
@@
III.6.2 Iterative Biconjugate Gradient Stabilized Method
@@
where $\mathfrak{L}$ is a positive semi-definite Hermitian operator. The BiCGStab algorithm allows us to iteratively converge to the steady-state solution without requiring explicit matrix inversion or full diagonalization, making it particularly suitable for large-scale dissipative quantum systems.
@@
$|r^{(k)}\rangle=\mathfrak{L}\,|\rho^{(k)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (S46)
@@
$\mathcal{K}_{m}\bigl{(}\mathfrak{L},|r^{(0)}\rangle\bigr{)}=\mathrm{span}\Bigl{\{}\,|r^{(0)}\rangle,\,\mathfrak{L}|r^{(0)}\rangle,\,\mathfrak{L}^{2}|r^{(0)}\rangle,\dots,\mathfrak{L}^{m-1}|r^{(0)}\rangle\Bigr{\}}.$ (S47)
@@
    $|r^{(k+1)}\rangle=\mathfrak{L}\,|\rho^{(k+1)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (S48)
@@
$\langle\hat{O}\rangle=\frac{\operatorname{Tr}(\hat{O}\hat{\rho}_{\mathrm{ss}})}{\operatorname{Tr}(\hat{\rho}_{\mathrm{ss}})},$ (S49)
@@
$\langle\hat{O}\rangle=\frac{\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\hat{O}|\hat{\rho}_{\mathrm{ss}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}{\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\mathds{I}|\hat{\rho}_{\mathrm{ss}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}.$ (S50)
@@
III.7 Evaluation of the Optimized Ansatz
After training our transformer density operator ansatz, we obtain an optimized parameter set $\bm{\theta}$ that approximates the steady-state density operator. To assess the accuracy of our transformer-based density operator ansatz, we compute the expectation values of local observables via Monte Carlo sampling, as discussed in Sec. III.1.

### v2

@@ C.2 Mixed-State Observables
$\langle\hat{A}\rangle=\frac{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\,\hat{A}\bigr{)}}{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\bigr{)}}=\sum_{\bm{\alpha}\in\mathbf{M}}\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\bigr{)}}\,\tilde{A}_{\rho_{\bm{\theta}}}(\bm{\alpha}),$ (43)
@@
$\tilde{A}_{\rho_{\bm{\theta}}}(\bm{\alpha})=\sum_{\bm{\beta}}\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}\langle\bm{\beta}|\hat{A}|\bm{\alpha}\rangle.$ (44)
@@
plays the role of a classical distribution over the diagonal elements of $\hat{\rho}_{\bm{\theta}}$, and the local estimator $\tilde{A}_{\rho_{\bm{\theta}}}(\bm{\alpha})$ here involves an inner sum over all basis states (or a suitably chosen subset) to capture the off-diagonal contributions $\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})$ to the observable $\hat{A}$. In this manner, the quantum expectation value reduces to a standard classical average over a series of sampled configurations $\mathbf{M}$, allowing one to use the same Metropolis-Hastings sampling scheme described in Sec. C.1 to estimate both observables and their gradients for the mixed-state variational ansatz.
C.3 Metric Tensor in Stochastic Reconfiguration
@@
$S_{ij}=\langle\Delta O_{i}^{*}\Delta O_{j}\rangle-\langle\Delta O_{i}^{*}\rangle\langle\Delta O_{j}\rangle,$ (45)
@@
$\Delta O_{i}=\frac{\partial\log\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}{\partial\theta_{i}}$ (46)
@@
In the context of mixed-state variational ansatz, this metric tensor is computed using Monte Carlo sampling as described in Sec. C.2. Once $S$ is constructed, it is used in the natural gradient update step to precondition the gradient of the energy functional.
C.4 Regularization and Stabilization
@@
$S^{\prime}=S+\lambda\mathds{I},$ (47)
@@
This is implemented as the stochastic reconfiguration method as a gradient preconditioner in NetKet, where $S$ is constructed from Monte Carlo estimates as described in Section C.3. This approach stabilizes the training dynamics of the transformer density operator ansatz and improves convergence in the steady-state optimization.
C.5 Optimization Procedure
@@
^$\hat{\mathcal{L}}|\rho_{\bm{\theta}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\hat{\mathcal{L}}\bm{\rho}_{\bm{\theta}}=0,$ (48)
@@
$\operatorname{Loss}(\bm{\theta})=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho_{\bm{\theta}}|\mathfrak{L}|\rho_{\bm{\theta}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (49)
@@
$\mathfrak{L}=\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}.$ (50)
@@
$\operatorname{Cost}(\bm{\theta})=\frac{\|\hat{\mathcal{L}}\bm{\rho}_{\bm{\theta}}\|_{2}^{2}}{\|\bm{\rho}_{\bm{\theta}}\|_{2}^{2}}=\frac{\operatorname{Tr}\Bigl{(}\bm{\rho}_{\bm{\theta}}^{\dagger}\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}\bm{\rho}_{\bm{\theta}}\Bigr{)}}{\operatorname{Tr}\Bigl{(}\bm{\rho}_{\bm{\theta}}^{\dagger}{\bm{\rho}_{\bm{\theta}}}\Bigr{)}},$ (51)
@@
In our implementation, we use a hybrid optimization approach that combines standard first-order gradient updates with second-order corrections via stochastic reconfiguration introduced in C.4. The Stochastic Gradient Descent or Adam optimizer performs standard gradient updates, while the stochastic reconfiguration accounts for the curvature of the variational manifold by introducing a metric tensor $S$, effectively implementing a natural gradient descent strategy. Their methods are provided in NetKet via a dedicated variational driver nk.SteadyState and Eq. (51) is the cost function that NetKet uses in its steady-state driver.
C.5.1 Hybrid Optimization Approach
@@
${\bm{\theta}}^{(k+1)}={\bm{\theta}}^{(k)}-\eta\nabla_{\bm{\theta}}\operatorname{Cost}(\bm{\theta}),$ (52)
@@
${\bm{\theta}}^{(k+1)}={\bm{\theta}}^{(k)}-\eta S^{-1}\nabla_{\bm{\theta}}\operatorname{Cost}(\bm{\theta}),$ (53)
@@
The stochastic gradient $\nabla_{\bm{\theta}}\operatorname{Cost}(\bm{\theta})$ is estimated over the probability distribution defined by the entries of the vectorized density matrix $P_{\bm{\theta}}(\bm{\alpha},\bm{\beta})\propto|\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})|^{2}$ as in Sec.C.2. The gradient of the cost function with respect to the complex conjugate of the $i$th parameter can be expressed as
$\frac{\partial}{\partial\theta_{i}^{*}}\operatorname{Cost}(\bm{\theta})=\langle\tilde{\mathcal{L}}_{i}\nabla_{i}^{*}\tilde{\mathcal{L}}_{i}\rangle-\langle O_{i}^{*}\,\tilde{\mathcal{L}}^{2}\rangle,$ (54)
@@
~$\tilde{\mathcal{L}}(\bm{\alpha},\bm{\beta})=\frac{\sum_{\bm{\alpha}^{\prime},\bm{\beta}^{\prime}}\hat{\mathcal{L}}(\bm{\alpha},\bm{\beta};\bm{\alpha}^{\prime},\bm{\beta}^{\prime})\,\rho_{\bm{\theta}}(\bm{\alpha}^{\prime},\bm{\beta}^{\prime})}{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}.$ (55)
C.5.2 Optimization Workflow
@@
    Sample a batch of configurations from the current transformer density operator ansatz using a Markov chain Monte Carlo sampler introduced in Sec. C.1.
@@
    Compute the metric tensor $S$ for stochastic reconfiguration introduced in Sec. C.3.
@@
    Regularize the metric tensor with a small constant $\lambda$, i.e., $S^{\prime}=S+\lambda\,\mathds{I}$, to ensure numerical stability as introduced in Sec. C.4.
@@
C.5.3 Learning Rate Scheduling
@@
$\mathrm{lr}(i_{\mathrm{step}})=\begin{cases}\eta_{0},&i_{\mathrm{step}}<i_{\mathrm{switch}}\\ \eta_{0}\cdot\frac{1+\cos(\pi(i_{\mathrm{step}}-i_{\mathrm{switch}})/i_{\mathrm{decay}})}{2}+\alpha\eta_{0},&i_{\mathrm{step}}\geq i_{\mathrm{switch}}\end{cases}$ (56)
@@
$\lambda_{\mathrm{SR}}(i_{\mathrm{step}})=\begin{cases}\lambda_{0},&i_{\mathrm{step}}<i_{\mathrm{switch,SR}}\\ \lambda_{0}\cdot\frac{1+\cos\left[\pi\left(i_{\mathrm{step}}-i_{\mathrm{switch,SR}}\right)/i_{\mathrm{decay,SR}}\right]}{2}+\alpha_{\mathrm{SR}}\lambda_{0},&i_{\mathrm{step}}\geq i_{\mathrm{switch,SR}}\end{cases}$ (57)
@@
C.6 Benchmark Methods for Steady-State Computation
@@ C.6.1 Exact Diagonalization
$\mathfrak{L}=\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}.$ (59)
@@
$\mathfrak{L}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\lambda|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (60)
@@
C.6.2 Iterative Biconjugate Gradient Stabilized Method
@@
where $\mathfrak{L}$ is a positive semi-definite Hermitian operator. The BiCGStab algorithm allows us to iteratively converge to the steady-state solution without requiring explicit matrix inversion or full diagonalization, making it particularly suitable for large-scale dissipative quantum systems. For instance, computing the steady state of a 13-site system using this iterative method requires substantial computational resources and time, typically taking several days even with optimized implementations, highlighting the computational advantages of our transformer-based approach.
@@
$|r^{(k)}\rangle=\mathfrak{L}\,|\rho^{(k)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (63)
@@
$\mathcal{K}_{m}\bigl{(}\mathfrak{L},|r^{(0)}\rangle\bigr{)}=\mathrm{span}\Bigl{\{}\,|r^{(0)}\rangle,\,\mathfrak{L}|r^{(0)}\rangle,\,\mathfrak{L}^{2}|r^{(0)}\rangle,\dots,\mathfrak{L}^{m-1}|r^{(0)}\rangle\Bigr{\}}.$ (64)
@@
    $|r^{(k+1)}\rangle=\mathfrak{L}\,|\rho^{(k+1)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (65)
@@
$\langle\hat{O}\rangle=\frac{\operatorname{Tr}(\hat{O}\hat{\rho}_{\mathrm{ss}})}{\operatorname{Tr}(\hat{\rho}_{\mathrm{ss}})},$ (66)
@@
$\langle\hat{O}\rangle=\frac{\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\hat{O}|\hat{\rho}_{\mathrm{ss}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}{\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\mathds{I}|\hat{\rho}_{\mathrm{ss}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}.$ (67)
@@
C.7 Evaluation of the Optimized Ansatz
After training our transformer density operator ansatz, we obtain an optimized parameter set $\bm{\theta}$ that approximates the steady-state density operator. To assess the accuracy of our transformer-based density operator ansatz, we compute the expectation values of local observables via Monte Carlo sampling, as discussed in Sec. C.1.